# Mingkai Liao

### v1

@@ 5.5 Real Datasets: Extended Settings
Figure 16: Real datasets (extended): AUC heatmaps across datasets (lower is better).
(a) ME-continuous.
(b) node_removal.
Figure 17: Real datasets (extended): AUC heatmaps across datasets (lower is better).
Discussion (heatmap). The heatmaps suggest PACIFIER’s gains are broadly distributed across datasets rather than driven by a small subset, while any occasional regressions tend to be localized to a few datasets, indicating robust cross-dataset performance.
Trajectory curves (representative; three datasets).
To visualize the timing of polarization reduction, we show representative trajectories on three datasets (chosen to cover small/medium/large graph sizes). Better methods reduce polarization earlier and keep it lower throughout the horizon. The complete set of per-dataset trajectory plots is provided in the Appendix.
Figure 18: Real datasets (MI-continuous): representative polarization trajectories on three datasets (lower and earlier is better). Full trajectories are deferred to the Appendix.
Figure 19: Real datasets (MI-bias_assmilate): representative polarization trajectories on three datasets (lower and earlier is better). Full trajectories are deferred to the Appendix.
Figure 20: Real datasets (ME-continuous): representative polarization trajectories on three datasets (lower and earlier is better). Full trajectories are deferred to the Appendix.
Figure 21: Real datasets (node_removal): representative polarization trajectories on three datasets (lower and earlier is better). Full trajectories are deferred to the Appendix.
5.5.1 Discussion
We next analyze the results under the extended FJ-consistent settings. For each variant (MI-continuous, MI-bias_assmilate, ME-continuous, and node_removal), we examine the empirical behavior from three complementary perspectives: (i) per-dataset AUC comparisons, (ii) cross-dataset heatmap patterns, and (iii) representative polarization trajectories that reveal the temporal dynamics of intervention. We discuss each setting individually and then summarize the broader robustness patterns across these extended dynamics.
MI-continuous.
Under continuous-valued opinion dynamics, the performance landscape changes substantially compared with the discrete MI setting. From the per-dataset AUC bars (Fig. 12), BOMP achieves the lowest average AUC (0.104), while PACIFIER-RL ranks second overall (avg. 0.114) and PACIFIER-Greedy follows closely behind. Across all 15 datasets, PACIFIER-RL outperforms the best non-PACIFIER baseline on 5 datasets (33% win rate), whereas PACIFIER-Greedy does not achieve wins against the strongest baseline.
The heatmap (Fig. 16(a)) reveals a clear size-dependent pattern. On small and mid-sized graphs (up to 15k nodes), BOMP remains consistently dominant, and both PACIFIER variants show a mild but systematic gap. However, on large graphs (>15,000>15{,}000 nodes), PACIFIER-RL becomes competitive and wins on 5 out of 6 datasets, with a slight positive average improvement (imp_mean $\approx$ 0.2%). This suggests that when graph scale increases and analytical structure becomes harder to exploit precisely, the learned policy begins to match or slightly exceed classical influence-based heuristics.
The representative trajectory curves (Fig. 18) further clarify this behavior. On small datasets (e.g., follow_germanwings), BOMP reduces polarization slightly earlier and maintains the lowest curve. On medium-scale datasets, PACIFIER-RL closely tracks BOMP but does not surpass it. On the largest dataset (retweet_mothersday), PACIFIER-RL achieves the best final AUC and maintains a marginally lower trajectory than all baselines, while PACIFIER-Greedy remains slightly behind.
Overall, MI-continuous exhibits a regime where analytical linear structure still provides a strong advantage on small graphs, but PACIFIER-RL demonstrates increasing competitiveness as scale grows. The gap between RL and Greedy indicates that even under continuous opinions, long-horizon credit assignment provides moderate benefits, though not as decisively as in the ME-style tasks.
MI-bias_assimilate.
Under nonlinear biased-assimilation dynamics, the relative ordering of methods changes markedly compared with the continuous linear case. From the per-dataset AUC bars (Fig. 13), both PACIFIER-Greedy (avg. 0.126) and PACIFIER-RL (avg. 0.135) substantially outperform all classical baselines, with PageRank being the strongest non-PACIFIER competitor (avg. 0.161). In contrast, BOMP—designed for linear steady-state structure—degrades significantly under biased assimilation.
Per-dataset comparisons show that PACIFIER-Greedy wins on 14 out of 15 datasets (93% win rate) against the best non-PACIFIER baseline, achieving an average relative improvement of approximately 19%. PACIFIER-RL also performs strongly (73% win rate, mean improvement $\approx$ 11%), though slightly behind the Greedy variant in aggregate. The heatmap (Fig. 16(b)) confirms that these gains are broadly distributed across datasets rather than concentrated on a few outliers.
A scale-aware pattern also emerges. On mid-sized graphs (5k–15k nodes), both PACIFIER variants achieve a 100% win rate and large improvement margins (mean $\approx$ 22–24%). On large graphs (>15>15k nodes), PACIFIER-Greedy remains consistently dominant (6/6 wins), while PACIFIER-RL remains competitive but slightly less stable. This suggests that under nonlinear opinion dynamics, immediate polarization reduction signals align well with long-term objectives, making myopic value estimation particularly effective.
The representative trajectory plots (Fig. 19) further illustrate this regime. On small graphs (e.g., follow_germanwings), both PACIFIER variants reduce polarization earlier than PageRank. On medium graphs, PACIFIER-Greedy often attains the lowest curve. On large graphs (e.g., retweet_mothersday), both variants substantially outperform classical baselines, with clear early-stage and sustained advantages.
Overall, MI-bias_assimilate demonstrates a setting where learning-based policies decisively outperform analytical linear-structure methods. Interestingly, the Greedy variant slightly surpasses RL on average, indicating that under biased assimilation, short-horizon gains already capture much of the effective intervention signal.
ME-continuous.
Under continuous expressed-opinion intervention (ME-continuous), a clear separation emerges between long-horizon reinforcement learning and myopic estimation. From the per-dataset AUC bars (Fig. 14), PACIFIER-RL achieves the best overall performance with a substantial margin (avg. 0.041), while the strongest classical baseline, PageRank, remains significantly higher (avg. 0.063). In contrast, PACIFIER-Greedy (avg. 0.081) performs similarly to BOMP and does not exhibit consistent gains.
Per-dataset comparisons highlight the dominance of RL: PACIFIER-RL wins on all 15 datasets (100% win rate), with an average relative improvement of approximately 34% over the best non-PACIFIER baseline. The improvement is stable across all scale buckets: small, mid, and large graphs each show consistent 30–35% average gains. By contrast, PACIFIER-Greedy wins only once (1/15 datasets), and its average improvement is strongly negative, particularly on mid- and large-scale graphs.
The heatmap (Fig. 17(a)) confirms that RL’s advantage is not driven by a small subset of datasets, but is uniformly distributed across scales. This pattern indicates that under ME-style dynamics, sequential dependency and long-horizon interaction effects are critical: the benefit of one intervention depends heavily on subsequent equilibrium shifts, which cannot be captured by immediate reward regression alone.
The representative trajectories (Fig. 20) further illustrate this behavior. On small graphs, RL reduces polarization substantially earlier and maintains the lowest trajectory throughout the horizon. On medium and large graphs, RL continues to dominate both PageRank and BOMP, while Greedy often diverges upward after early steps, reflecting poor long-term credit assignment.
[…]

### v2

@@ 5.6.2 Cost Setting
Figure 13: Real datasets (cost-ME): per-dataset bar comparison (lower is better).
Discussion.
We next analyze the real-world results under the cost-aware setting. Unlike the main real-world benchmark in Sec. 5.5, here we report only per-dataset bar comparisons, so the discussion focuses on cross-dataset performance patterns visible from these bars, together with subset-based summary statistics when BOMP is only partially available. A particularly important addition in this setting is the cost baseline, which constructs the intervention sequence simply by sorting nodes in ascending order of intervention cost. Because this rule ignores both opinion information and graph structure, its performance directly reveals how much of the task can already be solved by pure low-cost preference alone.
Evaluation protocol note.
As in the other real-world experiments, all cost-aware results are evaluated on the same 15 polarized topic graphs. However, BOMP remains only partially available due to its computational cost on larger instances. For both cost-MI and cost-ME, BOMP is available on 10 of the 15 datasets. Therefore, whenever BOMP is discussed, its averages should be interpreted only on the corresponding BOMP-comparable subset, rather than as directly comparable to full-coverage averages of methods such as PACIFIER. Accordingly, in the task-specific discussion below, we distinguish (i) comparisons on the BOMP-comparable subset and (ii) comparisons on the full dataset collection excluding BOMP from the strongest-baseline reference.
Cost-MI.
The cost-MI results reveal a qualitatively different pattern from the standard MI setting. Once intervention cost is introduced, the newly added cost baseline becomes immediately dominant among the non-PACIFIER methods, and in fact emerges as the strongest baseline overall on average. On the 10-dataset BOMP-comparable subset, cost achieves the best average AUC (0.158), slightly ahead of PACIFIER-Greedy (0.158) and more clearly ahead of PACIFIER-RL (0.166), while BOMP drops far behind (0.257). Thus, unlike the original MI benchmark where BOMP was the strongest method whenever available, under cost-aware internal-opinion moderation the primary competitive pressure no longer comes from the analytical solver, but from the extremely simple strategy of always choosing the cheapest nodes first.
This is a meaningful result in itself. It shows that in cost-MI, intervention cost is not a secondary nuisance variable but a central part of the optimization landscape. A baseline that ignores opinions and topology entirely, and uses only the cost weights, is already very hard to beat. This suggests that once the objective is cost-weighted, the benefit of sophisticated influence reasoning can be substantially compressed, because a large part of the achievable advantage is already captured by preferring low-cost interventions.
Against this strong baseline, the two PACIFIER variants behave quite differently. PACIFIER-RL does not outperform the strongest non-PACIFIER baseline on any dataset: its win rate is 0/10 on the BOMP-comparable subset and 0/15 on the full dataset collection excluding BOMP. Its mean improvement relative to the strongest non-PACIFIER baseline is negative in both views. By contrast, PACIFIER-Greedy is much more competitive. Although its average improvement is extremely small in magnitude, it stays essentially tied with cost overall and in fact wins on most datasets: 8/10 on the BOMP-comparable subset and 13/15 on the full dataset collection excluding BOMP. Its average AUC on the full dataset collection (0.164312) is also marginally lower than that of cost (0.164414).
This distinction between the two PACIFIER variants is highly informative. It suggests that cost-MI is a regime where strong performance depends less on long-horizon planning and more on making highly accurate local trade-offs between immediate polarization reduction and intervention cost. In such a regime, a myopic learned ranking policy is naturally better matched to the task than a more heavily bootstrapped RL objective. Indeed, PACIFIER-Greedy can be interpreted as learning a refined cost-aware priority rule that stays very close to the strong cost heuristic while still extracting a small but consistent extra gain on most datasets. PACIFIER-RL, on the other hand, appears to overemphasize longer-horizon structure that is less useful in this particular cost-aware internal-opinion setting.
Another notable consequence is that the relative ordering of classical baselines changes substantially once costs are introduced. In the original MI benchmark, BOMP was the clear strongest reference whenever available. Here, however, BOMP is no longer competitive, and the simple cost ordering dominates not only Random and the opinion-based heuristics, but also the analytical baseline itself. This indicates that the cost-aware extension is not a minor perturbation of the original problem, but rather a substantially different regime in which the dominant signal shifts from pure influence structure to cost-effectiveness.
Taken together, the cost-MI results support a nuanced conclusion. They do not show a broad PACIFIER victory in the same sense as the main ME experiments. Instead, they show that the newly introduced cost baseline is already extremely strong, that PACIFIER-Greedy is able to match and slightly exceed it overall, and that long-horizon RL is not always the best bias when the task is dominated by immediate cost-effectiveness. This makes cost-MI an important counterpoint in the paper: it highlights that PACIFIER is not uniformly superior in every extended setting, and that the relative merits of greedy versus RL policies depend strongly on the structure of the intervention objective.
Cost-ME.
The picture changes sharply in cost-ME. Here, the cost baseline again emerges as the strongest non-PACIFIER competitor, but unlike in cost-MI, it is now decisively surpassed by PACIFIER-RL. On the 10-dataset BOMP-comparable subset, PACIFIER-RL achieves the best average AUC (0.0685), well ahead of cost (0.0917), PACIFIER-Greedy (0.1007), and all classical baselines. Even on the full 15-dataset comparison excluding BOMP, PACIFIER-RL remains clearly best with an average AUC of 0.0711, compared with 0.0979 for cost. Its win rate reaches 9/10 on the BOMP-comparable subset and 14/15 on the full dataset collection, with mean improvements of 21.97% and 24.88%, respectively, over the strongest non-PACIFIER baseline.
This is perhaps the clearest evidence in the cost-aware experiments that PACIFIER-RL learns something substantially beyond simple cost preference. If cost alone were the dominant signal in cost-ME, then the cost baseline would be difficult to beat, as it is in cost-MI. Instead, PACIFIER-RL achieves large and highly consistent improvements over that baseline, which implies that effective cost-aware expressed-opinion moderation requires jointly reasoning about intervention cost, opinion state, and graph structure over the entire sequence. In other words, in cost-ME it is not enough to intervene cheaply; one must intervene cheaply and in the right places and order.
The behavior of PACIFIER-Greedy further reinforces this interpretation. Unlike in cost-MI, where the greedy variant is nearly tied with cost, in cost-ME it performs substantially worse. It wins on only 1/10 datasets on the BOMP-comparable subset and 1/15 on the full dataset collection excluding BOMP, with clearly negative mean improvements relative to the strongest non-PACIFIER baseline. Thus, in cost-ME, a myopic ranking rule is not sufficient. The fact that PACIFIER-RL succeeds while PACIFIER-Greedy does not indicates that longer-horizon sequential dependencies are genuinely important in this task.
This RL-versus-Greedy contrast also helps clarify the task difference between cost-MI and cost-ME. In cost-MI, the main challenge appears to be selecting cheap interventions with strong immediate payoff, which favors simple cost ordering and myopic learned refinement. In cost-ME, by contrast, intervening on one node changes the future expressed-opinion landscape in a way that makes later decisions strongly path-dependent. That is exactly the type of setting where bootstrapped long-horizon value learning should help, and the strong empirical advantage of PACIFIER-RL is fully consistent with that intuition.
It is also notable that the classical opinion-aware baselines remain very weak in this regime. Both ExtremeExpressed and ExtremeNeighbours perform far worse than cost, PACIFIER-Greedy, and especially PACIFIER-RL. This indicates that once explicit cost constraints are imposed, naive extremeness-based heuristics can become severely misaligned with the true optimization objective. The challenge is no longer simply to find influential or extreme nodes, but to find interventions with the best cumulative cost-adjusted effect over the trajectory.
Overall, the cost-ME results provide a strong positive result for PACIFIER. They show that the framework does not merely remain usable after adding cost constraints, but that its RL variant can substantially outperform both classical baselines and the newly introduced cost-only heuristic. This makes cost-ME one of the strongest demonstrations in the paper of the value of learned sequential planning under more realistic intervention constraints.
Overall picture in the cost-aware setting.
Taken together, the cost-aware results reveal a clear task-dependent pattern. For cost-MI, the simple cost baseline is already extremely strong, showing that much of the objective can be captured by pure low-cost preference. In this regime, PACIFIER-Greedy is the better PACIFIER variant: it essentially matches and slightly exceeds the cost-only rule overall, whereas PACIFIER-RL does not obtain an advantage. For cost-ME, by contrast, pure cost preference is no longer enough. Although the cost baseline remains the strongest non-PACIFIER competitor, PACIFIER-RL decisively outperforms it across almost all datasets, while PACIFIER-Greedy falls behind. This contrast strongly supports the broader claim of the paper: the value of PACIFIER depends on the degree to which the intervention task requires genuine sequential planning. When the problem is dominated by immediate cost-effectiveness, a greedy policy can be sufficient or even preferable; when future interaction effects matter, as in cost-ME, long-horizon RL becomes clearly advantageous.
Summary of real-world results.
Taken together, the real-world results reveal a clear and task-dependent overall pattern across both the canonical and extended settings. Throughout all these experiments, PACIFIER transfers non-trivially from synthetic training graphs to heterogeneous real-world networks, indicating that it captures meaningful intervention structure rather than merely overfitting to synthetic instances. At the same time, the strength and form of this transfer vary substantially across regimes, and these differences are themselves highly informative about when learning—and especially long-horizon RL—is most valuable.
For the analytically structured MI family, including both the standard and continuous settings, the strongest analytical baseline remains difficult to surpass whenever it is computationally feasible. In particular, BOMP continues to dominate in standard MI and continuous-MI, showing that when the problem remains closely aligned with linear influence structure, analytical reasoning still provides a very strong inductive bias. Nevertheless, PACIFIER already becomes highly competitive once the comparison shifts to scalable non-PACIFIER baselines, and PACIFIER-RL consistently emerges as the strongest learned or non-analytical alternative in these regimes. This indicates that even where learning does not fully replace analytical structure, it still captures useful intervention patterns that transfer meaningfully to real graphs.
A much stronger picture emerges for the ME family. In standard ME and continuous-ME, both PACIFIER variants consistently outperform all non-PACIFIER baselines across the real-world benchmark, demonstrating robust transfer of the learned intervention strategy beyond the synthetic training distribution. However, the cost-aware setting reveals a sharper separation between the two learned variants. In cost-ME, PACIFIER-RL substantially outperforms not only all classical baselines but also the strong cost-only baseline, whereas PACIFIER-Greedy falls clearly behind. This is one of the clearest signs in the paper that long-horizon reasoning matters: once intervention quality depends on how current budget-aware choices shape future expressed-opinion trajectories, RL gains a decisive advantage over myopic ranking.
The contrast with cost-MI is equally revealing. There, the newly introduced cost baseline becomes extremely strong, and PACIFIER-Greedy remains essentially tied with it, even slightly ahead overall, while PACIFIER-RL does not obtain an advantage. This suggests that cost-MI is dominated more by immediate cost-effectiveness than by deep sequential dependencies, and therefore does not reward long-horizon planning in the same way as cost-ME. Rather than weakening the overall argument, this contrast strengthens it: the experiments show that RL is most beneficial precisely in those settings where future interaction effects are genuinely important, not simply everywhere by default.
Finally, node_removal provides the strongest evidence that explicit sequential planning is essential. Here, PACIFIER-RL consistently and substantially outperforms both classical baselines and PACIFIER-Greedy, indicating that effective intervention depends critically on modeling the long-term structural consequences of each action. Among all the tasks considered, this is the regime where the distinction between myopic ranking and true long-horizon planning becomes most pronounced, and where the value of the RL formulation is most clearly exposed.
Overall, the real-world results show that PACIFIER achieves meaningful transfer across a broad spectrum of moderation regimes, while also revealing a sharp and interpretable pattern in the relative value of its two variants. PACIFIER-RL is not merely competitive in general; it is especially powerful in settings where analytical structure is weaker or where genuinely sequential, path-dependent planning is required, such as cost-ME and node_removal. By contrast, when the task is dominated by immediate local trade-offs, as in cost-MI, the gap between RL and myopic ranking narrows substantially. This regime-dependent picture is itself a central result of the paper: it shows not only that PACIFIER transfers to real networks, but also when and why long-horizon graph reinforcement learning provides its greatest advantage.
5.7 Runtime in the Standard Setting
We further compare the empirical runtime of all methods on the real-world datasets under the standard setting. Figure 14 presents two complementary views: (i) a comparison restricted to the node range covered by BOMP, and (ii) a full-range comparison after excluding BOMP. Together, these results illustrate the substantial scalability gap between BOMP and the remaining methods.
(a) Runtime comparison including BOMP, restricted to the maximum graph size that BOMP can handle.
(b) Full-range runtime comparison after excluding BOMP.
Figure 14: Runtime comparison on real-world datasets under the standard setting. Left: all methods are compared within the node-size range where BOMP remains executable. Right: after removing BOMP, the full-range scaling behavior of the remaining methods is shown. Overall, BOMP exhibits a much steeper growth trend than the other approaches, whereas PACIFIER remains scalable on large graphs.
Discussion.
The runtime results reveal a clear scalability distinction between BOMP and the remaining methods. According to the complexity analysis of Matakos et al., BOMP has time complexity $O(kn^{2})$, where $n$ is the number of nodes and $k$ is the intervention budget. Its high computational cost mainly comes from dense matrix-based candidate scoring built on the influence matrix $Q=(L+I)^{-1}$. In each iteration, BOMP repeatedly evaluates candidate nodes using dense influence information, which leads to a quadratic dependence on graph size and makes the method increasingly expensive on large graphs. By contrast, the lighter baselines used here are computationally cheaper: ExtremeExpressed has complexity $O(kn)$ in the MI setting, ExtremeNeighbours has complexity $O(k(n+m))$, and PageRank has complexity $O((m+n)I+n\log n)$, where $m$ is the number of edges and $I$ is the number of power-iteration steps.
This complexity gap is also clearly reflected in Fig. 14(a). Within the node range where BOMP remains executable, its runtime grows dramatically with graph size and reaches about $1.8\times 10^{4}$ seconds at the largest supported graph. In contrast, the other methods remain close to the bottom of the plot throughout the same range, and even the slowest among them is still lower by several orders of magnitude. This visual gap is consistent with the theoretical distinction between BOMP’s quadratic dependence on $n$ and the much lighter sparse-graph-oriented cost of the remaining methods.
For PACIFIER, the deployment-time complexity can be characterized directly from its feed-forward inference pipeline. In the application phase, the encoder performs a constant number $K$ of GraphSAGE-style message-passing rounds, which costs $O(K(n+m))$ on a sparse graph. The decoder then scores all candidate nodes in parallel, contributing an additional $O(n)$ term. If the final action is obtained through ranking or batch selection over all node scores, this adds an $O(n\log n)$ step. Therefore, the overall application-time complexity of PACIFIER is $O(K(n+m)+n\log n)$, which reduces to approximately $O(n+m+n\log n)$ when $K$ is treated as a small constant. Strictly speaking, this is not pure linear complexity in the formal sense, but it is near-linear in sparse graphs and is far milder than the quadratic $O(kn^{2})$ behavior of BOMP.
This near-linear scaling trend is consistent with Fig. 14(b). After excluding BOMP, all remaining methods show much gentler growth across the full graph range. Among them, PACIFIER is slower than the lightest baselines, especially on the largest graph where its runtime is around 89 seconds, while the next slowest method remains around 13 seconds. However, despite this moderate overhead, PACIFIER follows the same overall near-linear sparse-graph scaling regime as the other non-BOMP methods, rather than exhibiting the steep superlinear explosion observed for BOMP. This additional cost is acceptable in practice given its stronger moderation performance in the more challenging settings. Importantly, PACIFIER does not rely on dense influence-matrix operations for each candidate node during deployment. Instead, it uses a feed-forward inference pipeline in which the encoder and decoder directly produce action scores from graph representations. These differences explain why its empirical runtime grows much more gently and remains practical even on the largest real-world graphs in our benchmark.
5.8 Synthetic Benchmark: Full-Information Comparison
To further examine the role of intermediate steady-state information, we conduct an additional comparison on synthetic graphs under a full-information (FI) setting. Unlike the standard one-shot planning regime adopted throughout the main paper, the FI setting allows an algorithm to access the recomputed settled-opinion state after each partial intervention step. In other words, FI methods are allowed to use intermediate steady-state expressed opinions during sequential selection, and therefore do not satisfy the one-shot planning constraint defined in Sec. 3.4.
The purpose of this experiment is not to propose a new deployment setting, but to clarify how much performance can be gained when intermediate settled-opinion information is explicitly exposed to the planner. This comparison helps distinguish the contribution of learned planning under the practical one-shot regime from the contribution of additional oracle-like state information.
The compared methods follow the naming convention introduced in Sec. 5.3: methods with the suffix -FI are their full-information variants, while methods without this suffix remain in the standard one-shot regime. Accordingly, this benchmark includes the FI variants of heuristic and learned methods, together with a strong oracle-style greedy replanning baseline. Concretely, we compare ExtremeExpressed-FI, ExtremeNeighbours-FI, PACIFIER-RL, PACIFIER-Greedy, PACIFIER-RL-FI, PACIFIER-Greedy-FI, and Greedy. Among these methods, all entries marked with the suffix -FI as well as Greedy explicitly use intermediate settled-opinion information, whereas PACIFIER-RL and PACIFIER-Greedy do not. Therefore, this experiment directly measures the gap between one-shot planning and sequential replanning with full access to recomputed intermediate steady-state information.
We evaluate these methods on synthetic graphs under three representative settings: MI, ME, and topology-altering node-removal. The reported metric is the same trajectory-level accumulated normalized polarization (ANP) used in the main experiments, where lower values indicate better overall intervention quality.
(a) MI on synthetic graphs under full-information comparison.
(b) ME on synthetic graphs under full-information comparison.
(c) node-removal on synthetic graphs under full-information comparison.
Figure 15: Synthetic full-information comparison across three representative settings: MI, ME, and node-removal. Methods with the suffix -FI are allowed to access recomputed intermediate settled-opinion states during sequential selection, while methods without this suffix remain in the standard one-shot regime unless otherwise noted. The Greedy baseline is an oracle-style full-information replanning method that performs exhaustive one-step trial, rollback, and reselection at each decision step. Lower ANP indicates better trajectory-level polarization moderation performance.
Discussion.
The performance comparison in Fig. 15 leads to three main observations. First, exposing intermediate settled-opinion states is indeed helpful, but it does not fundamentally change the overall picture. Across MI, ME, and node-removal, methods with access to full information generally benefit from the additional feedback, yet the gains are typically incremental rather than transformative. In particular, the PACIFIER family remains highly competitive even when compared against baselines that are allowed to replan using recomputed intermediate steady states.
Second, the relatively small gap between the standard one-shot PACIFIER variants and their FI counterparts suggests that the main source of performance does not come merely from access to oracle-like intermediate information. Rather, it comes from the quality of the learned intervention policy itself. This point is especially important because PACIFIER-RL and PACIFIER-Greedy operate under a strictly weaker information regime than all -FI methods and the exhaustive Greedy baseline. That they still remain among the strongest methods indicates that effective long-horizon planning can recover most of the obtainable benefit without requiring sequential recomputation of the settled state after every step.
Third, the three synthetic settings also reveal that the value of FI depends on the intervention type. For MI and ME, FI access provides additional adaptivity during sequential selection, but learned planning remains the dominant factor in achieving low ANP. For topology-altering node-removal, the problem becomes more challenging because each intervention changes not only the opinion configuration but also the underlying interaction structure. In this setting, purely myopic improvement based on refreshed intermediate states is not sufficient to guarantee the best trajectory-level outcome, which highlights the importance of planning beyond immediate one-step gain.
Overall, the synthetic full-information comparison supports a clear conclusion: intermediate steady-state information is useful, but it is not the primary reason PACIFIER performs well. The proposed learned policies already capture most of the achievable moderation benefit under the practical one-shot regime, while FI-style sequential replanning mainly provides limited additional improvement rather than altering the qualitative conclusion of the study. This strengthens the case for PACIFIER as a practical approach: strong moderation performance can be achieved without assuming access to recomputed intermediate settled-opinion states during deployment.
5.9 Runtime in the Full-Information Setting
In practice, intermediate settled-opinion states are typically unavailable during deployment, which is one of the main reasons why our standard setting does not rely on such information. Therefore, the purpose of this runtime comparison is not to characterize practical deployment cost under realistic access assumptions, but to show, in a simulation setting, how much time different FI-style methods require when intermediate converged opinion states are explicitly recomputed and used during sequential selection.
We further compare the empirical runtime of all methods under the full-information (FI) setting. Figure 16 presents two complementary views: (i) a runtime comparison including all methods, and (ii) a comparison after removing the Greedy baseline. Together, these results reveal the substantial computational overhead introduced by FI-style sequential replanning.
(a) Runtime comparison of all methods under the full-information setting.
(b) Runtime comparison under the full-information setting after removing Greedy.
Figure 16: Runtime comparison on synthetic graphs under the full-information setting. Left: all methods are included. Right: the Greedy baseline is removed to make the differences among the remaining methods easier to observe. Overall, FI-based sequential replanning introduces substantial extra computational cost, while the standard PACIFIER remains markedly more efficient than the FI baselines.
Discussion.
The runtime results in Fig. 16 show a clear efficiency gap between the standard PACIFIER and FI-based methods when intermediate converged opinion states are explicitly recomputed in simulation.
From Fig. 16(a), the most striking pattern is the explosive growth of Greedy. As the average graph size increases from 40 to 450 nodes, its running time rises from nearly negligible values to about $1.25\times 10^{5}$, far exceeding all other methods. This shows that exhaustive full-information replanning becomes extremely costly once the graph size grows.
After removing Greedy, Fig. 16(b) makes the comparison among the remaining methods much clearer. The overall pattern is highly consistent: methods that rely on FI-style sequential replanning all incur substantially higher runtime, whereas the standard PACIFIER remains much faster across the full graph range. At 250 nodes, PACIFIER requires only about 75 runtime units, while the FI methods are already around 225–230 for ExtremeExpressed-FI and ExtremeNeighbours-FI, and above 400 for PACIFIER-FI. At 350 nodes, the standard PACIFIER is still only about 135, whereas the FI methods have increased to roughly 500 and above. At 450 nodes, PACIFIER remains around 210, while the three FI methods all stay much higher, ending around 920, 950, and 1410, respectively.
These results indicate that once intermediate converged opinion states are inserted into the decision loop, the runtime overhead rises sharply, regardless of the specific FI method used. In this sense, PACIFIER-FI, like the other FI algorithms, also bears the substantial computational burden brought by sequential full-information replanning. By contrast, the standard one-shot PACIFIER maintains a far more favorable runtime profile throughout all graph ranges.
Overall, this comparison supports a clear conclusion: although FI-style replanning can provide additional intermediate information in simulation, it comes with a pronounced computational cost. The standard one-shot PACIFIER, while maintaining strong performance, also demonstrates excellent computational efficiency. This advantage becomes increasingly clear as the graph size grows, further highlighting the practical value of the one-shot PACIFIER design.
5.10 Ablation Study
To validate the necessity of the two key representation components in PACIFIER, namely the polarization-aware auxiliary features and the temporal-aware node marking, we conduct a comprehensive ablation study on the real-world benchmark under both the MI and ME settings, each with two decision mechanisms: the greedy strategy and reinforcement learning. For each mechanism, we compare the full PACIFIER model against three ablated variants: removing the polarization-aware auxiliary features (without_aux), removing the first temporal marking component (without_mask1), and removing the second temporal marking component (without_mask2). The reported metric is ANP, where lower values indicate better trajectory-level moderation performance. For each dataset column, the best (lowest) score is highlighted in bold. The last column reports the average relative improvement of each ablated variant with respect to the corresponding full model.
Table 5: Ablation study under the MI setting with the greedy mechanism on real-world datasets (ANP; lower is better).
ANP follow_germanwings follow_mothersday retweet_wcw retweet_onedirection retweet_mothersday retweet_russia_march retweet_leadersdebate %
PACIFIER-Greedy 0.1359 0.2033 0.4101 0.2488 0.4000 0.3562 0.2372
PACIFIER-Greedy-without_aux 0.1532 (-12.71%) 0.2134 (-4.99%) 0.3978 (+3.01%) 0.2424 (+2.56%) 0.3943 (+1.42%) 0.3456 (+2.99%) 0.2396 (-1.01%) -0.62%
PACIFIER-Greedy-without_mask1 0.1330 (+2.15%) 0.2053 (-1.02%) 0.3982 (+2.90%) 0.2456 (+1.28%) 0.4047 (-1.17%) 0.3442 (+3.38%) 0.2362 (+0.39%) +1.45%
PACIFIER-Greedy-without_mask2 0.2383 (-75.33%) 0.2977 (-46.43%) 0.4704 (-14.70%) 0.3240 (-30.21%) 0.4506 (-12.65%) 0.4264 (-19.70%) 0.3246 (-36.88%) -29.05%
ANP retweet_jurassicworld retweet_germanwings retweet_nepal retweet_nationalkissingday retweet_ff retweet_gunsense retweet_ultralive retweet_sxsw
PACIFIER-Greedy 0.3633 0.3304 0.3200 0.1733 0.3894 0.3162 0.2124 0.3647
PACIFIER-Greedy-without_aux 0.3563 (+1.92%) 0.3285 (+0.58%) 0.3145 (+1.71%) 0.1819 (-4.97%) 0.3757 (+3.50%) 0.3111 (+1.61%) 0.2229 (-4.95%) 0.3642 (+0.13%)
PACIFIER-Greedy-without_mask1 0.3546 (+2.39%) 0.3245 (+1.81%) 0.3111 (+2.76%) 0.1734 (-0.06%) 0.3724 (+4.37%) 0.3078 (+2.67%) 0.2107 (+0.79%) 0.3678 (-0.86%)
PACIFIER-Greedy-without_mask2 0.4260 (-17.24%) 0.4002 (-21.13%) 0.3825 (-19.55%) 0.2448 (-41.26%) 0.4542 (-16.64%) 0.3904 (-23.47%) 0.3001 (-41.29%) 0.4351 (-19.31%)
Table 6: Ablation study under the MI setting with RL on real-world datasets (ANP; lower is better).
ANP follow_germanwings follow_mothersday retweet_wcw retweet_onedirection retweet_mothersday retweet_russia_march retweet_leadersdebate %
PACIFIER-RL 0.1328 0.2002 0.4087 0.2360 0.3980 0.3436 0.2285
PACIFIER-RL-without_aux 0.1573 (-18.43%) 0.2243 (-12.01%) 0.4010 (+1.89%) 0.2574 (-9.06%) 0.4041 (-1.53%) 0.3545 (-3.16%) 0.2551 (-11.64%) -5.38%
PACIFIER-RL-without_mask1 0.1341 (-0.98%) 0.2124 (-6.09%) 0.4253 (-4.06%) 0.2446 (-3.64%) 0.4036 (-1.40%) 0.3570 (-3.90%) 0.2354 (-3.02%) -3.30%
PACIFIER-RL-without_mask2 0.1572 (-18.34%) 0.2077 (-3.75%) 0.4071 (+0.39%) 0.2475 (-4.87%) 0.4027 (-1.18%) 0.3601 (-4.81%) 0.2471 (-8.14%) -4.36%
ANP retweet_jurassicworld retweet_germanwings retweet_nepal retweet_nationalkissingday retweet_ff retweet_gunsense retweet_ultralive retweet_sxsw
PACIFIER-RL 0.3612 0.3281 0.3033 0.1739 0.3827 0.3067 0.2064 0.3653
PACIFIER-RL-without_aux 0.3631 (-0.53%) 0.3423 (-4.33%) 0.3198 (-5.44%) 0.1760 (-1.21%) 0.3769 (+1.52%) 0.3204 (-4.47%) 0.2315 (-12.16%) 0.3658 (-0.14%)
PACIFIER-RL-without_mask1 0.3689 (-2.13%) 0.3292 (-0.34%) 0.3139 (-3.49%) 0.1876 (-7.88%) 0.4055 (-5.96%) 0.3095 (-0.91%) 0.2127 (-3.05%) 0.3749 (-2.63%)
PACIFIER-RL-without_mask2 0.3663 (-1.41%) 0.3397 (-3.53%) 0.3166 (-4.39%) 0.1732 (+0.40%) 0.3819 (+0.21%) 0.3202 (-4.40%) 0.2266 (-9.79%) 0.3720 (-1.83%)
Table 7: Ablation study under the ME setting with the greedy mechanism on real-world datasets (ANP; lower is better).
ANP follow_germanwings follow_mothersday retweet_wcw retweet_onedirection retweet_mothersday retweet_russia_march retweet_leadersdebate %
PACIFIER-Greedy 0.0619 0.0559 0.1187 0.0864 0.1159 0.1010 0.0888
PACIFIER-Greedy-without_aux 0.0597 (+3.48%) 0.0552 (+1.24%) 0.1213 (-2.17%) 0.0861 (+0.36%) 0.1188 (-2.51%) 0.1026 (-1.52%) 0.0902 (-1.63%) -0.93%
[…]